1 0
Add a comment Improve this question Transcribed image text
Answer #1

Let the orizinal length be L

let it is cut into a and L-a length

Original pipe f = v/2L

After cut

f1 = v/2*a

2a/v = 1/f1 ---------------(1)

f2 = v/2(L-a)

1/f2 = 2(L-a)/v = 2L/v - 2a/v

From eq(1) , put the value , we get

1/f2 = 1/f - 1/f1

f = f1f2/(f1 + f2)

= (256*440)/(256+440)

= 161.839 Hz

===================================

Therefore

v/2L = 161.839

343/(2L) = 161.839

L = 1.0596 m
Answer #2

L1=\lambda/2

L1=\lambda/2

frequency=Vsound/2*L1=256Hz

L1=Vsound/512

L2=\lambda/2

L2=\lambda/2

frequency=Vsound/2*L2=440Hz

L2=vsound/880

total length L=L1+L2

resonant frequency of total pipe=Vsound/2L

=Vsound/2*(L1+L2)

=Vsound/2*(Vsound/512+Vsound/880)

=162.003Hz

total length=Velocity of sound/324.006m